# DEM: Distribution Edited Model for Training with Mixed Data Distributions

### v1

@@ Appendix A Training Hyperparameters
Appendix B Choosing DEM Weights
In order to select the optimal mixing weights for DEM (Eq 2), we perform a grid search over $\omega_{i}$ values. For each coefficient combination we evaluate the validation losses on the five datasets used for fine-tuning (Section 5.1), and select the model that minimizes their average. However, exhaustive grid search is expensive as the number of combinations grows exponentially. Thus, we simplify Eq 2 and optimize a single coefficient $\omega$ for all datasets. We found $\omega=0.25$ (out of 10 values) to produce the best validation loss and use it for all our experiments.
@@ Appendix B Choosing DEM Weights
Appendix C Choosing Data Mixing Weights
Appendix D Fine-Grained Results

### v2

@@ Appendix A Training Hyperparameters
Appendix B Choosing Data Mixing Weights
Appendix C Choosing DEM Weights
@@
In order to select the optimal mixing weights for DEM - Distribution Vector (Eq 2), we perform a grid search over $\omega_{i}$ values. For each coefficient combination we evaluate the validation losses on the five datasets used for fine-tuning (Section 5.1), and select the model that minimizes their average. However, exhaustive grid search is expensive as the number of combinations grows exponentially. Thus, we simplify Eq 2 and optimize a single coefficient $\omega$ for all datasets. We found $\omega=0.25$ (out of 10 values) to produce the best validation loss and use it for all our experiments.
We chose the weights for DEM - Interpolation (Eq 3) in a similar manner as DEM - Distribution Vector by randomly sampling weights from the search grid and normalizing them to sum to 1. Additionally, we also tried the same weights as data mixing and equal weight of 0.2 for each of 5 datasets. The simplest strategy of equal weight performed on par with the best weight combination in terms of average val loss. So, we chose this and reported the corresponding results in Table 2